OpenStudy (anonymous):

find all real solutions of sq root of 5-x = x+7 (step by step response please) :)

OpenStudy (anonymous):

Is it \[\sqrt{5-x} = x+7\]or\[\sqrt{5}-x=x+7\]Use parentheses please

OpenStudy (anonymous):

(5-x)

OpenStudy (precal):

square both sides

OpenStudy (precal):

5-x=(x+7)^2

OpenStudy (precal):

5-x=x^2+14x+49 add x to both sides 5=x^2+15x+49 -5 to both sides 0=x^2+15x+44

OpenStudy (precal):

0=(x+11)(x+4) x=-11 x=-4

OpenStudy (anonymous):

thank you!

OpenStudy (precal):

anytime